Obtain the Fourier cosine series for the following function:


F(x)= 1 for (0 less than equal to x< 1)

0 for (1 less than equal to x < 4 )
1
Expert's answer
2019-07-19T11:12:23-0400

Formula for Fourier cosine series is

"F(x) \\approx a_{0} +\\sum_{n=1}^{ \t\\infty}a_{n}cos(\\frac{n*\\pi}{L}x)"

where

"a_{0} = \frac{1}{2L} \int_{-L}^{L}f(x) dx"

and

"a_{n} = \frac{1}{L} \int_{-L}^{L}f(x)cos(\frac{n\pi x}{L}) dx"


if function is even, then

"\int_{-L}^{L}f(x) dx = 2\int_{0}^{L}f(x) dx"

"a_{0} = \frac{1}{4} \int_{0}^{1}1 dx = \frac{1}{4}"


"a_{n} = \frac{1}{2}\int_{0}^{1}cos(\frac{n \pi x}{4}) dx=\frac{2sin(\frac{\pi n}{4})}{\pi n}"

so Fourier cosine series is


"F(x) = \frac{1}{4} +\sum_{n=1}^{ \infty}\frac{2sin(\frac{\pi n}{4})}{\pi n}"